# Quantum Polymorphisms and Commutativity Gadgets

### v1

@@ Abstract.
We introduce the concept of quantum polymorphisms to the complexity theory of non-local games. We use this notion to give a full characterisation of the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions. Prior to our work, a classification was only known in the Boolean case [Culf–Mastel, STOC’25]. As an application of our framework, we prove that the entangled CSP parameterised by odd cycles is undecidable. Furthermore, we establish a quantum version of Galois connection for entangled CSPs in the case of non-oracular quantum homomorphisms.
1. Introduction
In 1971, Cook’s introduction of $\mathsf{NP}$-completeness [cook1971complexity] provided a way to mathematically assess the inherent hardness of computational problems. Boolean satisfiability, graph colouring, linear equations over finite fields, and many other computatational problems of a seemingly distinct nature systematically began to fall into either the polynomial-time solvable or the $\mathsf{NP}$-complete buckets [Karp72], while intermediate complexity turned out to be a rarer phenomenon [ladner1975structure]. This fact soon sparked the hypothesis that the broad, natural class of computational problems involving satisfaction of constraints over finite alphabets—collectively known as (finite-domain) CSPs and encompassing the aforementioned problems and many others—could never exhibit the unnatural intermediate behaviour. $\operatorname{P}$ vs. $\mathsf{NP}$-complete’ dichotomies for fragments of CSPs started to emerge, primary examples being Schaefer’s [Schaefer78:stoc] and Hell–Nešetřil’s [HellN90] dichotomies for Boolean and graph CSPs, respectively. However, a unified framework for approaching the CSP Dichotomy Conjecture [Feder98:monotone] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [Jeavons97:closure, Jeavons98:algebraic]. By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [Bulatov17:focs] and, independently, Zhuk [Zhuk17_FOCS, Zhuk20:jacm] in 2017.
Seven years before the publication of Cook’s result, the work of Bell [bell1964einstein] originating from the Einstein–Podolsky–Rosen paradox [einstein1935can] demonstrated that quantum systems can display correlations unattainable within classical physics, thus ruling out local hidden-variable theories as realistic explanations for quantum mechanics. In its mathematical essence, Bell’s result can be formulated as a property of CSPs: There exist specific CSP languages for which the classical value and the entangled value can differ, where the latter is defined as the optimal success probability in a multiplayer interactive proof of the satisfiability of the CSP instance where the players have access to quantum resources. A primary question arising from Bell’s theorem concerns the computational complexity of determining the entangled value of a CSP. In particular, it is known that determining whether an instance of SAT is perfectly quantum satisfiable is undecidable (in fact, even the gap version of the problem is known to be undecidable, as a consequence of the celebrated $\operatorname{MIP^{*}=RE}$ theorem [ji2021mip]). On the other hand, examples of tractable CSP languages for which quantum and classical satisfiability coincide, which implies polynomial-time solvability of the corresponding entangled CSPs, are known [atserias_generalized_2019, ciardo_quantum_minion, bulatov2025satisfiability].
@@ 1. Introduction
Nevertheless, unlike in the classical case, the complexity landscape of entangled CSPs remains largely unexplored. A natural strategy to shed some light on it is to transfer the algebraic machinery that allowed proving the (classical) CSP Dichotomy to the quantum setting. In particular, for classical CSPs, the concept of polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level. The root of this phenomenon is a Galois connection between relational clones (expressing the CSP language) and operation clones (expressing the family of polymorphisms). Via this connection, interreducibility between CSPs associated with different languages can be explained via inclusion (or weaker forms of the latter, in particular, so called height-1 clone homomorphisms) between their polymorphism clones.
@@
To what extent can this machinery by transferred to the setting of entangled CSPs, and thus provide some understanding of their complexity and interreducibility? This question was first addressed at this level of generality by Ji in [ji2013binary]. In particular, that work identified a substantial obstacle against the prospect of quantising all types of algebraic CSP reductions known as gadgets. Ultimately, the obstacle is caused by the non-classical phenomenon of contextuality: Replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability. In the same paper, Ji provided a method to solve this issue for specific CSP languages, via the new notion of commutativity gadgets. On a high level, these are specific types of gadgets having no classical counterpart, that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction. For example, by showing that a specific triangular prism construction yields a commutativity gadgets for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding entangled CSPs, thus proving undecidability of quantum 3-colouring.

### v2

# Quantum Polymorphisms and the Complexity of Quantum Constraint Satisfaction

@@ Abstract
We introduce the concept of quantum polymorphisms to the complexity theory of quantum constraint satisfaction. Via this notion, we build an algebraic framework of reductions between quantum CSPs, and we establish a Galois connection between quantum polymorphism minions and quantum relational constructions. By leveraging a contextuality property of quantum polymorphisms, we fully characterise the existence of commutativity gadgets for relational structures, introduced by Ji as a method for achieving quantum soundness of classical CSP reductions. Prior to our work, only a partial classification was known for a subclass of Boolean languages and for non-Boolean languages meeting specific structural conditions [Culf–Mastel, FOCS’25]. As an application of our framework, we prove that the quantum CSPs parameterised by odd cycles and the quantum CSP expressing quantum satisfiability of Siggers clauses are undecidable.
1 Introduction
The quantum constraint satisfaction problem is the decision problem of testing for the existence of a winning quantum strategy for a 2-player 1-round game where the two (cooperating, computationally unbounded) players aim to prove to a classical verifier that a system of constraints can be simultaneously satisfied. Here, a quantum strategy consists of a bipartite finite-dimensional quantum state shared by the players and a collection of measurements that the players perform on the state. Upon receiving from the verifier questions involving local constraints of the system, the players measure their part of the state and answer with partial assignments based on the measurement outcomes. The origins of such game—often known as a non-local game—come from physics: Bell’s[12] and Kochen–Specker’s [53] proofs of the incompatibility of quantum mechanics with local hidden-variable theories, stemming from the Einstein–Podolsky–Rosen paradox [35], depend on the fact that, for certain types of constraints, quantum strategies generate correlations in the player’s answers that are unattainable in classical physics.
[…]
However, a unified framework for approaching the CSP Dichotomy Conjecture [36] only emerged at the turn of the century, with the foundational line of work initiated by Jeavons, Cohen, and Gyssens that became known as the (universal-)algebraic approach to CSPs [44, 45]. By lifting CSP theory from the realm of Turing machines and reductions to that of algebra, the study of polymorphism clones—universal-algebraic objects capturing high-order invariance properties of CSP solutions—provided a clearer viewpoint on the CSP complexity landscape, and it eventually culminated with the positive resolution of the conjecture by Bulatov [21] and, independently, Zhuk [68, 69] in 2017. In the case of quantum CSPs, the complexity landscape is much less understood. While there are various examples of constraint classes for which the quantum CSP is undecidable [67, 46, 41, 2, 33, 63, 34] or polynomial-time solvable [50, 47, 2, 30, 63, 20, 56], the complexity is unknown for most classes of languages, due to the lack of a general approach for studying them with unified tools.
The main conceptual contribution of this work is the introduction of the notion of quantum polymorphisms as a tool to lift the study of algebraic invariants of classical CSP languages to the quantum setting. Via this notion, we initiate an algebraic complexity theory of quantum CSPs modelled on the algebraic approach initiated in [44, 45] and culminated 20 years later with the proof of the (classical) Dichotomy Theorem.
@@ 1 Introduction
Classical polymorphisms turned out to be the right tool for exploring CSP complexity at the algebraic level. The root of this phenomenon is a connection between relational constructions (expressing the CSP language) and operation systems (expressing the family of polymorphisms). Via this connection, reducibility between CSPs associated with different languages can be explained via inclusion (or algebraic forms of the latter, in particular, so called minion homomorphisms) between their polymorphism sets. In this work, we prove that quantum polymorphisms completely determine the complexity of quantum CSPs, thus enabling the use of algebra in the complexity analysis of such problems. Moreover, we link quantum polymorphisms to a quantised version of classical relational constructions that we name q-definitions and q-constructions, obtained by lifting the classical model-theoretic counterparts to the setting of quantum strategies.
@@
Broadly, computable reductions in the literature on non-local games and quantum CSP complexity typically follow a common conceptual pattern: taking a classical CSP reduction (in particular, a so-called subdivision of constraints or a more general pp-definition), and trying to achieve quantum completeness and soundness. It is well known that this approach does not work in general. A substantial obstacle to the prospect of using classical gadgets as reductions between quantum CSPs is ultimately caused by the following, non-classical phenomenon: replacing constraints by gadgets increases the size of the contexts—which, in general, destroys simultaneous measurability (see, e.g., [46, 63, 2]). On the other hand, for some families of constraint languages, the issue can be solved by pairing classical gadgets with an extra type of gadget known as a commutativity gadget. On a high level, these are specific types of gadgets that can be plugged into classical gadgets to enforce simultaneous measurability of the final instance resulting from the reduction. For example, by showing that a specific triangular prism construction yields a commutativity gadget for the 3-clique, Ji showed that the classical polynomial-time reduction from 3-SAT to 3-colouring can be lifted to a reduction between the corresponding quantum CSPs [47]. These commutativity gadgets have no counterpart in the classical setting, as they enforce (by design) no additional constraints on the variables in their scope apart from commutativity. Therefore, they are useless from the classical point of view.